# Investigating Disordered Granular Matter via Ordered Geometric Fragmentation

### v1

@@ V Summary
Fragmentation is implemented through a well-defined slicing algorithm, applied iteratively and described in detail in the main text. At each stage of fragmentation, the resulting fragments are reassembled into configurations of maximal admissible volume, referred to as towers. The volumes of these towers were analysed as functions of the fragmentation stage. In realistic granular systems composed of elongated grains, the constituents at a given stage of fragmentation are expected to arrange in a largely unconstrained manner, resulting in occupied volumes that lie between those of compact packing and the largest geometrically admissible configurations. The construction of maximal-volume towers therefore provides a purely geometric upper bound and allows the characteristic scale and qualitative trends of volume changes induced by fragmentation to be estimated.
@@
It has been shown that, at the initial stage of fragmentation, the occupied volume increases relative to that of the parent object. With further fragmentation, the volume then decreases monotonically, approaching at the final stage the universal value $5/4$ of the original volume. It follows that, throughout the fragmentation process, a material composed of elongated grains typically occupies a volume larger than that of the original object, unless the grains are packed perfectly without voids. Within the fragmentation sequence, pairs of configurations have been identified and termed conjugate towers. These towers are composed of geometrically indistinguishable building blocks but enclose different volumes. When the number of fragmentation stages is even, a self-conjugate tower appears at the centre of the sequence. The magnitude of the volume difference between conjugate towers depends on their separation along the fragmentation sequence: towers corresponding to nearby stages exhibit smaller differences, whereas the largest difference occurs between the first and final stages of fragmentation.
Explicit relations have also been derived, and a procedure has been outlined for connecting the idealised geometric model to systems of real cuboidal grains and for experimentally testing its predictions. These include the monotonic evolution of volume with fragmentation and the emergence of phase-like transitions between distinct packing configurations driven purely by geometric rearrangement.
@@
In summary, an analytically tractable geometric model of fragmentation and reassembly has been presented, revealing universal features of volume evolution and phase-like behaviour in granular systems. Despite its simplicity, the model captures essential geometric aspects of granular matter and provides a useful framework for the investigation of packing geometry, porosity, and metastable configurations.
@@ Acknowledgements
This work is dedicated to the memory of the author’s grandmother, Liza. Many years ago, she remarked that a given mass of corn kernels occupies a smaller volume than the flour obtained from them. This simple observation provided the original motivation for the present study and inspired the investigation of the geometric origins underlying such volume changes.

### v2

@@ V Summary
Fragmentation is implemented through a well-defined slicing algorithm, applied iteratively and described in detail in the main text. At each stage of fragmentation, the resulting fragments are reassembled into configurations of maximal admissible volume, referred to as towers. The volumes of these towers were analysed as functions of the fragmentation stage. In realistic granular systems composed of elongated grains, the constituents at a given stage of fragmentation are expected to arrange in a largely unconstrained manner, resulting in occupied volumes that lie between those of compact packing and the largest geometrically admissible configurations. The construction of maximal-volume towers therefore provides a purely geometric upper bound on occupied volume, or equivalently, a lower bound on packing fraction, and allows the characteristic scale and qualitative trends of volume changes induced by fragmentation to be estimated.
@@
It has been shown that, at the initial stage of fragmentation, the occupied volume increases relative to that of the parent object. With further fragmentation, the volume then decreases monotonically, approaching at the final stage the limiting value $5/4$ of the original volume. It follows that, throughout the fragmentation process, a material composed of elongated grains typically occupies a volume larger than that of the original object, unless the grains are packed perfectly without voids.
The inverse of the relative volume, corresponding to the packing fraction, was shown to obey the asymptotic scaling law $\phi\propto 1/\alpha$ for large aspect ratios, in agreement with experimental observations. Comparison with experimental data from Freeman et al. [9] demonstrates that the model provides a geometric lower bound on packing fraction, with experimental values lying consistently above the theoretical predictions, as expected for realistic random packings.
Within the fragmentation sequence, pairs of configurations have been identified and termed conjugate towers. These towers are composed of geometrically indistinguishable building blocks but enclose different volumes. When the number of fragmentation stages is even, a self-conjugate tower appears at the centre of the sequence. The magnitude of the volume difference between conjugate towers depends on their separation along the fragmentation sequence: towers corresponding to nearby stages exhibit smaller differences, whereas the largest difference occurs between the first and final stages of fragmentation.
Transitions between conjugate towers may be interpreted as geometric phase transitions. A criterion for the observability of such transitions has been derived, showing that they are restricted to small systems satisfying $N_{g}\lesssim 4l_{g}/a$. In large bulk systems where this criterion is not globally satisfied, geometric phase transitions may nevertheless occur locally within small domains of aligned grains. X-ray tomography studies of rod packings reveal such local alignment patterns, with characteristic domain sizes of order $N_{\text{domain}}\sim 4l_{g}/a$ grains, consistent with the theoretical prediction. A directly testable experimental signature is proposed: domain size should scale linearly with the aspect ratio $\alpha$, providing a quantitative test of the geometric phase transition mechanism.
@@
In summary, an analytically tractable geometric model of fragmentation and reassembly has been presented, revealing features of volume evolution and phase-like behaviour in granular systems. Despite its simplicity, the model captures essential geometric aspects of granular matter, provides quantitative predictions in agreement with experimental scaling laws, and offers a framework for investigating packing geometry, porosity, metastable configurations, and local ordering phenomena in elongated grain assemblies.
@@ Acknowledgements
The author is grateful to A. Zaccone for his interest in this work and for suggesting the review article Ref. [4].
The author also thanks E. R. Weeks for kindly providing numerical data from the experiments reported in Ref. [9].
This work is dedicated to the memory of the author’s grandmother, Liza Tsertsvadze. Many years ago, she remarked that a given mass of corn kernels occupies a smaller volume than the flour obtained from them. This simple observation provided the original motivation for the present study and inspired the investigation of the geometric origins underlying such volume changes.
The author acknowledges the use of AI language models (ChatGPT and Claude) to assist with manuscript preparation, including text editing, LaTeX formatting, and literature organization. All scientific content, theoretical development, and conclusions are solely the work of the author.